# Multi-Agent Collaborative Data Selection for Efficient LLM Pretraining

### v1

@@ 3.4 Multi-agent Collaboration
1:Training data $\mathcal{D}$, reference task $\mathcal{D}_{\text{ref}}$, main model $\mathcal{M}$, optimizer $\mathcal{O}$, total training steps $T$, selected size $k$, update step $U$, Memory for each agent $\mathcal{H}_{\mathcal{A}}$
@@
6:         for each agent $\mathcal{A}$ do
7:              Sample data points uniformly from agent search space
@@
9:              Update agent weight $\mathbf{w_{\mathcal{A}}}\leftarrow\mathbf{w_{\mathcal{A}}}+\eta_{\mathcal{A}}\cdot\mathbf{\overline{R}_{\mathcal{A}}}$
@@
11:         Compute agent score $\overline{R}_{\mathcal{A}}$ and average score $\overline{R}$ according to Eq. 8
@@
13:         Calculate coordinator score $S(x_{i})$ for $x_{i}\in\mathcal{D}$ according to Eq. 7
@@
Ultimately, the agent console defined in Definition 2 aggregates all agents’ feedback to compute a final score for each data point, determining the final data selection decision.
Multi-agent collaboration.
In the context of multi-agent collaboration, the weighted score for each agent must be calculated to evaluate their respective contributions effectively. This calculation takes into account various factors specific to each agent. For every data sample $x_{i}$, the overall score $S(x_{i})$ is determined by the following formula:
$S(x_{i})=(\theta_{\text{Quality}}\cdot\mathcal{S}_{\text{Quality}}(x_{i})+\theta_{\text{Domain}}\cdot\mathcal{S}_{\text{Domain}}(x_{i})+\theta_{\text{Topic}}\cdot\mathcal{S}_{\text{Topic}}(x_{i})),$ (7)
where $\mathcal{S}_{\text{Quality}}(x_{i})$, $\mathcal{S}_{\text{Domain}}(x_{i})$, and $\mathcal{S}_{\text{Topic}}(x_{i})$ are the scores calculated by the quality, domain, and topic agents for the sample $x_{i}$, respectively. And $\mathbf{\theta_{\mathcal{A}}}\in\{\theta_{\text{Quality}},\theta_{\text{Domain}},\theta_{\text{Topic}}\}$ is the collaborative weight for each agent, which is updated during training process.
Collaborative weight update.
To dynamically adjust the importance of each agent during various training phases, we modify the agent’s collaborative weight based on its overall rewards. We compute the reward of each agent and the average reward across all agents:
$\overline{R}_{\mathcal{A}}=\frac{1}{|n|}\sum_{j=1}^{n}w_{\mathcal{A}}^{j}\cdot\overline{R}_{\mathcal{A}}^{j},\quad\overline{R}=\frac{1}{k}\sum\overline{R}_{\mathcal{A}},$ (8)
This information is then used to update each agent’s collaborative weight, which is stored in the agent console’s memory for future decision-making:
By continuously refining these weights, the collaboration strategy adapts to optimize overall performance and appropriately adjust the role of each agent throughout different stages of training. The complete training pipeline is outlined in Algorithm 1.
4 Experiments
We conduct a series of experiments to evaluate the effectiveness of our multi-agent collaborative data selection method. Comprehensively, we find that: (1) In the end-to-end experiments, our approach introduces significant improvement in terms of data efficiency leading to faster convergence for LLM training, and achieves up to $10.5\%$ improvements on average across various language model benchmarks when compared with other baseline approaches (§4.1); (2) We also verify that the design and implementation of the core components in our multi-agent framework design are necessary to reach this advanced performance through a set of carefully designed ablation studies (§4.2).
Table 1: Our approach improves model performance across multiple tasks. To ensure all demonstrations fit within the 1024-token context window, we present a comprehensive results table for 0-shot, 3-shot, and 5-shot scenarios in Appendix A.3. The selected tasks in this table include: Problem Solving: ARC-Easy (3), ARC-Challenge (3), MathQA (3), MMLU (3); Commonsense Reasoning: OpenBookQA (3), SocialIQA (3), Winogrande (0), CommonsenseQA (5); and Reading Comprehension: BoolQ (0), Race (5). For the QuRating and DSIR methods, we use their best-performing variants: QuRating-Edu and DSIR-Wiki, respectively.
Selection Method Problem Solving Commonsense Reasoning Reading    Comprehension Average (4 tasks) (4 tasks) (2 tasks) (10 tasks) Random sampling - 30B tokens 31.1 32.9 43.1 34.2 Random sampling - 60B tokens 33.6$\uparrow$2.5 33.7$\uparrow$0.8 46.1$\uparrow$3.0 36.1$\uparrow$1.9 Perplexity PPL (Ankner et al., 2024) 29.9$\downarrow$1.2 30.5$\downarrow$2.4 42.4$\downarrow$0.7 32.7$\downarrow$1.5 Classifier-based data selection QuRating (Wettig et al., 2024) 34.1$\uparrow$3.0 34.1$\uparrow$1.2 41.4$\downarrow$1.7 35.6$\uparrow$1.4 FineWeb-Edu (Penedo et al., 2024) 32.6$\uparrow$1.5 33.0$\uparrow$0.1 45.3$\uparrow$2.2 35.3$\uparrow$1.1 DSIR (Xie et al., 2023b) 30.9$\downarrow$0.2 32.0$\downarrow$0.8 41.5$\downarrow$1.6 33.5$\downarrow$0.7 Domain mixing methods DOGE Fan et al. (2024) 30.9$\downarrow$0.2 32.2$\downarrow$0.7 45.1$\uparrow$2.0 34.3$\uparrow$0.1 DoReMi (Xie et al., 2023a) 30.4$\downarrow$0.7 32.6$\downarrow$0.3 44.8$\uparrow$1.7 34.1$\downarrow$0.1 DMLaw (Ye et al., 2024) 30.2$\downarrow$0.9 32.1$\downarrow$0.9 45.1$\uparrow$2.0 33.9$\downarrow$0.3 RegMix (Liu et al., 2024) 30.7$\downarrow$0.4 32.5$\downarrow$0.4 44.6$\uparrow$1.5 34.2$\uparrow$0.0 Influence MATES (Yu et al., 2024) 30.9$\downarrow$0.2 34.0$\uparrow$1.1 46.5$\uparrow$3.4 35.3$\uparrow$1.1 Multi-agent collaboration (ours) 36.7$\uparrow$5.6 34.8$\uparrow$1.9 45.9$\uparrow$2.8 37.8$\uparrow$3.6
4.1 End-to-end Experiments
We evaluate our multi-agent framework against a wide category of state-of-the-art approaches to compare the data efficiency for LLM pretraining. We train a $1.3$ billion parameter Llama-2 architecture model with $30$ billion selected tokens.
  • Pretraining datasets. We utilize the popular SlimPajama (Soboleva et al., 2023) dataset including $670$ billion tokens, which is derived from the RedPajama (Computer, 2023) dataset. The SlimPajama (Soboleva et al., 2023) provide the meta-data about the domain information for each sample. Before the training process, we annotate the entire dataset using the FineWeb-Edu quality scorer (Penedo et al., 2024) along with our custom-trained BERT-based topic classifier. The training details for the topic classifier is provided in Appendix §A.1.
Discussion. We highlight that our proposed multi-agent collaborative data selection mechanism introduces statistical efficiency in terms of LLM training convergence and also provides some computational efficiency in terms of data processing overheads. In terms of statistical efficiency, our method consistently outperforms others at every benchmarked training step, as shown in Figure 4. While MATES Yu et al. (2024) performs comparably to our methods during the early training phase (steps $1500$ to $3000$), its performance drops in later stages. This aligns with its original paper, which notes that relying solely on influence functions for specific reference tasks (e.g., LAMBADA (Paperno et al., 2016)) can degrade performance in mid-to-late pretraining. Despite this, MATES still outperforms other methods without dynamic adjustments shown in Figure 4. In contrast, our multi-agent collaborative data selection mechanism can dynamically adjust the corresponding weights from different agents and select data based on the most up-to-date model preferences, effectively mitigating biases and surpassing other domain-mixing and data-selection techniques. In terms of computational efficiency, we also achieve higher computational efficiency than previous methods. For example, QuRating Wettig et al. (2024) requires around $2000$ NVIDIA A800 GPU hours to label the entire SlimPajama dataset, while our offline labeling takes just $180$ GPU hours. MATES Yu et al. (2024), which recalculates influence scores and trains a BERT model for each labeling cycle, incurs $360$ GPU hours for a four-stage update. Additionally, MATES’ labels are only usable in the next training stage, making it time-consuming and difficult to scale. In contrast, our method can improve the computational efficiency from two aspects: (1) we find that a group of light-weight agents collaboratively enables superior data selection, which is more computational efficiency than any method that requires a heavy data processing or label procedure; (2) the collaborative, dynamic learning procedure introduced in our multi-agent framework is computational efficient; by using a sampled holdout set and CPU-based calculations for updating agent parameters, our computational overhead is ignorable compared with heavy LLM training computation.
4.2 Ablation Study
We introduce a set of carefully designed ablation studies to justify the design and implementation of our multi-agent collaborative data selection framework. Concretely, (1) we test the combination of different agents to show the advance introduced by collaboration, and (2) we verify the necessity of the dynamic adjustment of the agent’s weight for data selection.
Experiment setups. The dataset used in our experiments is identical to that in the end-to-end experiment. We perform our ablation study using a 373M model. The complete model architecture is detailed in Appendix §A.2-Table 3. We train all these $373$M models using 8 A800 GPUs for $5000$ steps. Our evaluation tasks are similar to those in the end-to-end results, but we specifically report the in-context learning performance for 3-shot evaluation.
@@ 4.2 Ablation Study
Table 2: This ablation study examines the performance of various combinations of agent collaboration and update mechanisms. Accuracy is reported for all tasks, with the highest value in each column shown in bold.
@@
Problem Solving Commonsense Reasoning Reading Compreh. Selection Method ARC-E ARC-C MathQA MMLU O.B.QA SIQA W.G. C.S.QA BoolQ RACE Average Quality&Domain&Topic Agent 57.9 24.7 21.9 25.4 20.2 37.9 52.6 20.4 59.6 29.4 35.0 without collaboration update 47.9 20.4 21.0 25.1 17.2 37.3 51.3 20.0 56.5 28.3 32.5 Domain&Quality Agent 55.1 18.6 21.7 24.4 17.4 37.1 51.2 19.8 61.7 28.2 33.5 Topic&Quality Agent 56.2 24.4 21.8 25.2 19.4 36.3 49.0 19.7 56.1 28.5 33.6 Domain&Topic Agent 44.6 18.3 21.7 25.7 16.2 36.6 51.9 19.9 61.6 27.8 32.4 Quality Agent 53.0 24.7 21.8 25.5 18.0 36.3 49.5 18.1 57.0 28.0 32.9 Domain Agent 44.1 19.1 20.8 25.6 16.6 36.8 52.0 19.7 56.7 28.2 32.0 Topic Agent 42.7 19.2 21.0 27.0 17.4 37.1 50.7 19.7 54.6 28.5 31.8 No Agent 42.5 20.0 21.1 23.8 14.6 35.9 50.1 18.8 56.1 27.9 31.1
@@
Results and discussion. The results of the ablation study are shown in Table 2. We want to highlight the result from two aspects: First, the ablation study underscores the importance of each agent in achieving optimal performance across the training tasks. When the quality, domain, and topic agents are used together, the model performs best, highlighting the benefits of their combined use, as shown in Table 2. In terms of evaluating the performance of each agent, we find that the quality agent outperforms other single-agent configurations, excelling in problem solving tasks. However, its performance drops by $7.6\%$ on tasks requiring domain knowledge or contextual understanding. Here, the domain and topic agents play a crucial role, as they excel in these areas. Despite this, neither performs well on problem solving tasks, except for the topic agent, which significantly improves MMLU performance, indicating that topic diversity may benefit such tasks. In terms of evaluating the combination of the agents, we find that removing any agent noticeably reduces overall accuracy, though the impact varies. Excluding the quality agent leads to the largest drop, over $7.4\%$, significantly affecting performance in problem solving tasks, and commonsense reasoning tasks like OpenbookQA. This highlights the quality agent’s vital role in reasoning and problem-solving. Similarly, excluding the topic agent causes a performance drop in ARC-Challenge and a $4\%$ reduction in MMLU, emphasizing its importance in tasks covering diverse subjects; removing the domain agent results in a $5\%$ performance drop in commonsense reasoning tasks, underscoring its key contribution to these areas. Second, the ablation study verifies the design and implementation of the collaborative dynamic adjustment of the agents’ weights (introduced §3.4) for efficient data selection. Concretely, in this variant, all agents were initialized with equal weights, which remained fixed throughout training without adjusting for individual agent performance. Surprisingly, this fixed-weight approach (equal to random sampling) resulted in a significant $7.1\%$ drop in overall performance compared to the dynamic weighting used in the collaborative update framework, as shown in Table 2. We believe this result from the ablation study is a strong indicator that the dynamic adjustment of the celebration mechanism is essential for efficient data selection.
5 Related Work
Data selection in LLM pretraining. Selecting high-quality pretraining data from large corpora is crucial for effective LLM training. Recent approaches leverage various methodologies for efficient data selection. Concretely, classifiers (Brown, 2020; Chowdhery et al., 2023; Du et al., 2022; Xie et al., 2023b) and language modeling perplexity (Wenzek et al., 2020; Thrush et al., 2024) have been applied to identify data resembling high-quality samples; recently, more advanced quality scores based on classifier have shown the effectiveness in data selection, e.g., QuRating (Wettig et al., 2024), FineWeb-Edu (Lozhkov et al., 2024), etc. Data mixture is another effective way to improve data diversity, at both token level (Touvron et al., 2023a; Gao et al., 2020; Soboleva et al., 2023) and sample level, e.g., DoReMi (Xie et al., 2023a), DOGE (Fan et al., 2024), DMLaw (Ye et al., 2024), and RegMix (Liu et al., 2024); very recently, topic distributions has also been considered as an effective data mixing method, e.g., the downsampling overrepresented topics in Llama 3.1 (Team, 2024). Influence functions have been studied to understand for data efficiency  (Koh & Liang, 2017), and some recent attempts based on efficient approximation have been proposed to improve data efficiency in LLM pretraining (Schioppa et al., 2022; Grosse et al., 2023; Isonuma & Titov, 2024); for example, MATES (Yu et al., 2024) uses a staged BERT model to assess data influence, QUAD (Zhang et al., 2024) leverage cluster information to reduce the computational cost of calculating individual data influence.
Multi-agent collaborative frameworks. Multi-agent collaborative frameworks (Russell & Norvig, 2016; Wooldridge, 2009) facilitate cooperative problem solving among autonomous agents and have been widely applied to solve various problems, e.g., neural optimizer search (Bello et al., 2017), collebrative LLM programming (Hong et al., ). In these systems, agents may have conflicting goals and independently take actions based on their objectives; a reward mechanism evaluates these actions, providing feedback that allows each agent to adjust and refine its strategy over time; a central console coordinates the agents by synthesizing their feedback and guiding the overall system towards more optimal decisions (Russell & Norvig, 2016; Wooldridge, 2009). Collaboration among agents is dynamic, as they adapt their ability to work together improves (Olfati-Saber, 2006).
6 Conclusions
In this paper, we addressed the critical challenge of efficient data selection for LLM pretraining by proposing a novel multi-agent collaborative data selection mechanism. Our framework allows multiple data selection methods to operate as independent agents, with an agent console designed to dynamically integrate their outputs throughout the LLM training process. Through extensive empirical studies, we demonstrated that our approach significantly improves data efficiency, leading to faster convergence during LLM pretraining and achieving an average performance improvement up to $10.5\%$ across various language model benchmarks compared to state-of-the-art methods. Such promising results highlight the effectiveness of combining multiple data selection strategies dynamically, offering a solution to resolve the potential inherent conflicts from data selection methods to provide superior data efficiency for LLM pretraining.
Appendix A Appendix
A.1 Details of Training Topic Classifier
A.2 Details of Pretraining
A.2.1 Details of Pretraining Models Architecture

### v3

@@ A.4.2 Actor Design
1:Training data $\mathcal{D}$, reference task $\mathcal{D}_{\text{ref}}$, main model $\mathcal{M}$, optimizer $\mathcal{O}$, total training steps $T$, selected size $k$, update step $U$, Memory for each actor $\mathcal{H}_{\mathcal{A}}$
@@
6:         for each actor $\mathcal{A}$ do
7:              Sample data points according to actor’s predefined sampling distribution
@@
9:              Update actor weight $\mathbf{w_{\mathcal{A}}}\leftarrow\mathbf{w_{\mathcal{A}}}+\eta_{\mathcal{A}}\cdot\mathbf{\overline{R}_{\mathcal{A}}}$
@@
11:         Compute actor score $\overline{R}_{\mathcal{A}}$ and average score $\overline{R}$ according to Eq. 7
@@
13:         Calculate coordinator score $S(x_{i})$ for $x_{i}\in\mathcal{D}$ according to Eq. LABEL:eq:coordinator
@@
A.4.3 Multi-actor Collaboration
Assume that the score of a single data point in the reference task is obtained as a weighted sum of multiple components. The total score for each data point is given by Equation 7 as:
$S(x_{i})=\sum_{\mathcal{A}}\theta_{\mathcal{A}}\cdot S_{\mathcal{A}}(x_{i}),$
where $\theta_{\mathcal{A}}$ are collaborative weights. A central coordinator adjusts these weights over time based on the actors’ contributions to the overall reward:
$\theta_{\mathcal{A}}(t+1)=\theta_{\mathcal{A}}(t)+\eta_{\mathcal{A}}(\bar{R}_{\mathcal{A}}-\bar{R}),$
where $\bar{R}_{\mathcal{A}}$ is the actor’s average reward, and $\bar{R}$ is the global average reward:
$\bar{R}_{\mathcal{A}}=\frac{1}{n}\sum_{j=1}^{n}w_{\mathcal{A}}^{j}\cdot\bar{R}_{\mathcal{A}}^{j},\quad\bar{R}=\frac{1}{3}\sum_{\mathcal{A}}\bar{R}_{\mathcal{A}}.$
We consider three possible cases for our framework, comparing its relationship with traditional optimization problem.
  • Single-actor case: If only one actor is involved, $\theta$ becomes irrelevant, reducing the problem to a classical optimization scenario where the actor greedily selects the optimal data based on one criteria.
  • Multi-actor competitive mechanism: When multiple actors are present, $\theta$ reflects each actor’s capability. Selecting the best-performing actor for decision-making introduces a heuristic competitive mechanism, building upon the classical optimization framework.
  • Multi-actor collaborative mechanism: Alternatively, when multiple actors are involved, $\theta$ can be used to weigh each actor’s contributions for decision-making. This introduces a smoother heuristic cooperative mechanism, extending the classical optimization framework by leveraging weighted collaboration. This heuristic cooperative mechanism dynamically adjusts the influence of each actor based on the model’s current preferences, enabling more effective data filtering decisions.
In practice, we choose to use the multi-actor collaborative mechanism for data selection. We have added comparisons with single-actor and competitive mechanisms in Table 5 to further elaborate the effectiveness of collaboration.
@@ A.4.3 Multi-actor Collaboration
Table 5: This ablation study examines the performance of various combinations of actor collaboration (Actor) and dynamic collaborative weight update (Dynamic). Accuracy is reported for all tasks, with the highest value in each column shown in bold.
@@
Problem Solving Commonsense Reasoning Reading Compreh. actor Dynamic ARC-E ARC-C MathQA MMLU O.B.QA SIQA W.G. C.S.QA BoolQ RACE Average with with 65.8 31.5 23.0 26.6 24.6 39.9 54.1 20.1 60.4 30.5 37.7 with without 59.4 26.3 21.3 25.1 20.5 38.9 52.9 19.8 58.1 28.3 35.1 without - 59.2 26.1 20.3 25.4 21.3 39.1 52.6 20.1 56.5 29.1 35.0
A.5 Detailed Analysis of Computational Overhead
In this subsection, we compare the computational overhead of our multi-actor collaboration framework with baseline approaches. The analysis focuses on three aspects: offline computation efficiency, online computation efficiency, and overall FLOPs requirements. Table 6 summarizes these comparisons.
Table 6: Comparison of Computational Overhead Across Methods
Selection Method Offline Computation Online Computation Overall
Cost (FLOPs) Cost (FLOPs) (FLOPs)
Qu-Rating (Wettig et al., 2024) $7.13\times 10^{20}$ N.A. $7.13\times 10^{20}$
MATES (Yu et al., 2024) N.A. $1.99\times 10^{20}$ $1.99\times 10^{20}$
Multi-actor collaboration (ours) $9.91\times 10^{19}$ $1.19\times 10^{18}$ $1.00\times 10^{20}$
A.5.1 Offline Computation Efficiency
Our method achieves superior offline efficiency by requiring only $9.91\times 10^{19}$ FLOPs for a one-time dataset labeling process using a 109M BERT-based model for inference. This is nearly an order of magnitude more efficient than Qu-Rating, which consumes $7.13\times 10^{20}$ FLOPs due to its reliance on a larger 1.3B Sheared-LLaMA model. MATES does not utilize offline computation, relying solely on online updates, which avoids this cost but limits its flexibility and scalability. The offline labeling in our method ensures robust initial scores for large-scale datasets while laying the groundwork for efficient online updates.
A.5.2 Online Computation Efficiency
For adaptive online updates, both our approach and MATES compute influence scores with $1.19\times 10^{18}$ FLOPs. However, MATES involves labeling the entire dataset with a 109M BERT-based model in every round, amounting to $1.98\times 10^{20}$ FLOPs across four data selection stages. In contrast, our method avoids re-labeling the entire dataset, significantly reducing the computational cost by focusing on labeling the large pretraining datasets only once.
Overall, our approach cuts the computational cost in half compared to MATES and requires only about 1/7 of the computational resources used by Qu-Rating.
A.6 Analysis of Ablation Study
A.6.1 Analysis of actor Roles on Different Type of Tasks
Table 7: This ablation study examines the performance of various combinations of actor collaboration and update mechanisms. All models are in 373M LLaMA2 architecture. Accuracy is reported for all tasks, with the highest value in each column shown in bold.
Problem Solving Commonsense Reasoning Reading Compreh. Selection Method ARC-E ARC-C MathQA MMLU O.B.QA SIQA W.G. C.S.QA BoolQ RACE Average Quality&Domain&Topic actor 57.9 24.7 21.9 25.4 20.2 37.9 52.6 20.4 59.6 29.4 35.0 without collaboration update 47.9 20.4 21.0 25.1 17.2 37.3 51.3 20.0 56.5 28.3 32.5 Domain&Quality actor 55.1 18.6 21.7 24.4 17.4 37.1 51.2 19.8 61.7 28.2 33.5 Topic&Quality actor 56.2 24.4 21.8 25.2 19.4 36.3 49.0 19.7 56.1 28.5 33.6 Domain&Topic actor 44.6 18.3 21.7 25.7 16.2 36.6 51.9 19.9 61.6 27.8 32.4 Quality actor 53.0 24.7 21.8 25.5 18.0 36.3 49.5 18.1 57.0 28.0 32.9 Domain actor 44.1 19.1 20.8 25.6 16.6 36.8 52.0 19.7 56.7 28.2 32.0 Topic actor 42.7 19.2 21.0 27.0 17.4 37.1 50.7 19.7 54.6 28.5 31.8 No actor 42.5 20.0 21.1 23.8 14.6 35.9 50.1 18.8 56.1 27.9 31.1
We show the actor ablation study conducted on 373M LLaMA2 models Table  7 as well as 1.3B LLaMA2 models Table 2.
@@ A.6.1 Analysis of actor Roles on Different Type of Tasks
First, the ablation study underscores the importance of each actor in achieving optimal performance across the training tasks. When the quality, domain, and topic actors are used together, the model performs best, highlighting the benefits of their combined use, as shown in Table 2. In terms of evaluating the performance of each actor, we find that the quality actor outperforms other single-actor configurations, excelling in problem solving tasks. However, its performance drops on tasks requiring domain knowledge or contextual understanding. Here, the domain and topic actors play a crucial role, as they excel in these areas. Despite this, neither performs well on problem solving tasks, except for the topic actor, which significantly improves MMLU performance, indicating that topic diversity may benefit such tasks. In terms of evaluating the combination of the actors, we find that removing any actor noticeably reduces overall accuracy, though the impact varies. Excluding the quality actor leads to the largest drop, significantly affecting performance in problem solving tasks, and commonsense reasoning tasks like OpenbookQA. This highlights the quality actor’s vital role in reasoning and problem-solving. Similarly, excluding the topic actor causes a performance drop in ARC-Challenge and a significant reduction in MMLU, emphasizing its importance in tasks covering diverse subjects; removing the domain actor results in a performance drop in commonsense reasoning tasks, underscoring its key contribution to these areas. Second, the ablation study verifies the design and implementation of the collaborative dynamic adjustment of the actors’ weights (introduced §3.4) for efficient data selection. Concretely, in this variant, all actors were initialized with equal weights, which remained fixed throughout training without adjusting for individual actor performance. Surprisingly, this fixed-weight approach (equal to random sampling) resulted in a significant drop in overall performance compared to the dynamic weighting used in the collaborative update framework, as shown in Table 2. We believe this result from the ablation study is a strong indicator that the dynamic adjustment of the celebration mechanism is essential for efficient data selection.
A.6.2 Ablation Study on Reference Tasks Selection
Table 8: This ablation study examines the performance of various combinations of reference task. Accuracy is reported for all tasks, with the highest value in each column shown in bold.
Problem Solving Commonsense Reasoning Reading Compreh. Reference Tasks ARC-E ARC-C MathQA MMLU O.B.QA SIQA W.G. C.S.QA BoolQ RACE Average LAMBADA&SQuAD&Jeopardy 65.8 31.5 23.0 26.6 24.6 39.9 54.1 20.1 60.4 30.5 37.7 LAMBADA 64.3 31.2 22.3 26.8 23.5 39.6 54.6 20.4 59.6 30.1 37.2 SQuAD 65.1 30.9 23.4 25.9 24.9 40.1 53.8 21.2 59.1 29.3 37.4 Jeopardy 63.9 30.3 23.6 26.3 24.1 40.7 54.5 21.8 59.1 30.2 37.5 Random selection 54.6 23.0 22.1 24.9 18.8 40.3 52.9 21.5 53.0 29.8 34.1
In our experiments of selecting reference tasks in Table 8, we observe that while the choice of reference tasks can influence performance, the impact on average performance is marginal (within 0.5 points). Using different reference tasks consistently leads to a significant improvement in average performance compared to random data selection, demonstrating that our method is not sensitive to the choice of reference tasks.
A.7 Details of Pretraining
A.7.1 Details of Pretraining Models Architecture